# Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning

### v3

@@ 1 Introduction
Continual Learning (CL) [45, 27] is a research direction that focuses on realizing the human dream of creating truly intelligent systems, where machines can learn on the fly, accumulate knowledge, and operate in constantly changing environments as a human’s companion. Despite the impressive capabilities of A.I systems, continual learning remains a challenging scenario due to the tendency to forget obtained knowledge when facing new ones, known as catastrophic forgetting [12]. In dealing with this challenge, traditional CL methods often rely on storing past data for replaying during new tasks [28, 6], which can raise concerns about memory usage and privacy. Besides, prior work shows that replay methods result in overfitting and poor generalization [28, 42, 7]. To overcome these limitations, recent methods proposed leveraging the strong generalization ability of pretrained models [14, 19] to solve sequences of CL tasks. A notable line of work is the prompt-based approach [48, 38, 25], where a small set of learnable prompts is injected to pretrained backbones for adapting emerging tasks over time.
@@
In addition, we discover that human natural learning behavior has many valuable aspects, especially the habit of analyzing data, organizing them in a meaningful way, and finding connections between old and new knowledge [37, 3, 39, 30]. Inspired by these practices, we investigate the characteristics of current common benchmark datasets, as well as the behavior of pretrained models. Our findings reveal that the data can always be categorized into consistent groups, regardless of their arrival times. Each group usually includes classes with similar semantic characteristics that models may confuse and, thus, should be paid special attention to during training. Identifying these kinds of groups will benefit in improving model representation learning, as well as aiding in addressing catastrophic forgetting.
Therefore, we propose a training strategy that constantly considers emerging data in groups, following a hierarchical tree-like taxonomy developed based on expert knowledge. In particular, during the training on new tasks, the feature extractor is trained to distinguish all classes so far, especially focusing on classes within the same group. We observe that images belonging to concepts/labels within the same leaf group in the taxonomy share strong visual and semantic correlations, leading to overlapping representations in the latent space, which compromises performance. By encouraging the feature extractor to separate and contrast their representations more distinctly, we can reduce the overlap of easily confused classes, thereby further improving performance. This strategy not only mitigates forgetting when new classes emerge but also consolidates domain-specific knowledge in each of these leaf groups.
@@
  • Originating from the research findings of cognitive science, we propose a novel approach to reduce forgetting by exploiting relationships between data. By dynamically constructing label-based hierarchical taxonomies and leveraging the prior knowledge of pretrained models via an optimal transport-based approach, we can identify the challenging knowledge areas that need further focus during the sequence of tasks.
  • We empirically evaluate the effectiveness of our method against current state-of-the-art pretrained-based baselines across various continual learning benchmarks.
Organization. The rest of the paper is structured as follows. In Section 2, we present related work. Then in Section 3, we formulate the problem and introduce a new perspective to explain the cause of forgetting in prompt-based CL models. Section 4 transitions from the motivation provided by Cognitive Science insights to the proposed training strategy, emphasizing the importance of leveraging relationships between class data. Section 5 presents the experimental results, and finally, we discuss the limitations and suggest future directions in Section 6.
@@ 2 Related Work
Class Incremental Learning.
This is one of the most challenging and widely studied CL scenarios [41, 44], where task identity is unknown during testing, and data of previous data is inaccessible during current training [29, 36, 16, 13]. This work follows the setting of CIL and proposes a novel approach to mitigate forgetting and improve performance for prompt-based CL models.
@@
This line of work exploits the power of pretrained backbone to quickly adapt to the sequence of downstream tasks by updating just a small number of parameters or prompts. Initial work like [49, 48, 38] typically assign a set of prompts to tasks, enhancing the adaptability of the backbone to downstream tasks. However, the absence of explicit constraints can lead to feature overlapping between classes from different tasks. Therefore, recent methods employ some types of contrastive loss [44, 24] or utilize Vision Language models [46, 33] to better separate features from tasks. However, they treat all classes equally during training, missing the opportunity to learn in challenging areas where classes have many similarities and are easily confused. In this work, we propose a novel approach to exploit the relationships within data, allowing the model to recognize groups of these classes, and develop a deeper understanding of the respective knowledge areas, thereby reducing forgetting and enhancing its ability to learn new tasks.
@@ 3 Preliminary Analysis
In this section, we first present the problem formulation of prompt-based continual learning in Section 3.1. Following this, we analyze the sources of forgetting in these methods in Section 3.2.
@@ 3.1 Problem formulation
Dataset Sup-21K DINO-1K
Split-Imagenet-R 3.82 80.31
Split-CIFAR100 1.32 48.43
Split-CUB-200 2.04 63.56
(a) Distribution shift of representations corresponding to $\mathcal{D}_{1}$ after learning sequence of tasks.
(b) Within task accuracy on $\mathcal{D}_{1}$ (Sup-21K)
(c) True accuracy on $\mathcal{D}_{1}$ (Sup-21K)
Figure 1: Empirical study about forgetting of HiDe-Prompt. Table (a) reports L2 Wasserstein distance between the latent distributions corresponding to $\bm{P}_{t}$ and $\bm{P}_{\hat{t}}$ of data task 1, after learning the final task; Figure (b) Accuracy of the first task over time, using classification head $s(\bm{x})$; Figure (c) Accuracy of the first task over time, using classification head $h_{\psi}$.
@@ 3.2 Forgetting in Prompt-based Methods
In the context of Continual Learning, forgetting refers to the phenomenon in which performance on previously learned tasks decreases over time. Current prompt-based CL methods, which leverage the power of pretrained models, attribute forgetting either to (I) changes in parameters from the backbone when using the common prompt pool $\bm{P}$ for all tasks [49, 48] or to (II) the inherent mismatch between the models used at training and testing, as discussed and analyzed in [53, 40]. In this work, we complement these views by providing empirical studies to offer new insights about the reason for forgetting in this type of model, which arises from overlapping between old and new emerging class representations.
@@
To uncover the new factor causing forgetting, we propose to consider the cases where the behavior of backbone w.r.t each learned task is nearly unchanged over time - meaning (I) and (II) roughly would not happen. Firstly, to eliminate concerns about changing learned parameters (I), we consider methods that propose using a distinct set of prompts $\bm{P}_{t}$ to a specific task $t$. Then, the remaining potential factor is the difference between the prompt chosen at inference time, which we refer as $\bm{P}_{\hat{t}}$, and the one used during training $\bm{P}_{t}$. Thus, we conduct experiments on HiDe-Prompt [44], which is the latest SOTA in prompt-based CL, to measure the differences between the features formed when using these two prompts. Particularly, we consider $\bm{x}\in\mathcal{D}_{1}$ that belongs to the first task and measure the L2 Wasserstein distance $W_{2}(Q,\hat{Q})$ [20] between $Q$ (i.e., the latent distribution corresponding to $\bm{P}_{t}$, which consists of $f_{\Phi,\bm{P}_{{t}}}(\bm{x})$ for $\bm{x}\sim\mathcal{D}_{1}$) and $\hat{Q}$ (i.e., the latent distribution corresponding $\bm{P}_{\hat{t}}$ after learning the last task, consists of $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$ for for $\bm{x}\sim\mathcal{D}_{1}$). The results in Table 1(a) show that the difference between these distributions is negligible when using Sup-21K as a pre-trained backbone, in stark contrast to the significant shift observed in DINO-1K.
@@
For a detailed examination of Sup-21K, we propose an experimental setup to analyze the performance degradation of corresponding models over time. Specifically, we train a classifier $s(\bm{x})$ on the latent space $f_{\Phi,\bm{P}_{1}}(\bm{x})$ derived from the first task $\mathcal{D}_{1}$ to accurately classify the classes within this initial task. As the model undergoes continual training on a sequence of subsequent tasks, we assess the ability of $s(\bm{x})$ to classify instances from $\mathcal{D}_{1}$ based on the latent space representation $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$, defining this as within-task accuracy. This is compared to the true accuracy when using the classifier $h_{\psi}$. The results in Figure 1(b), show that within-task accuracy remains nearly stable, suggesting minimal shift between representations $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$ and $f_{\Phi,\bm{P}_{t}}(\bm{x})$. In contrast, Figure 1(c), illustrates a significant decrease in true accuracy, raising the question of whether we have overlooked additional factors contributing to final forgetfulness.
@@
Additionally, considering the inference feature space of $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$, we can see that as more tasks arrive, the number of classes increase, making the space fuller and increasing the possibility of overlap between class distributions. To support this point, we illustrate the representation of some class images in Figure 2, as well as the corresponding t-SNE visualizations in Figure 4. In particular, after Task 1, we have representations of ”oak tree”, ”mouse” and ”porcupine” located in quite separate locations. However, when Task 2 and then Task 3 arrive, the appearance of ”willow tree” and ”pine tree” makes the latent space become fuller, and ”oak tree” no longer maintains the separation from the remaining classes as before, leading to a remarkable drop in model performance.
@@
Figure 2: Problem: When new classes arrive, the latent space of a model for all tasks so far becomes fuller, and class representations tend to be overlapped, leading to performance degradation in old tasks. Our solution: In this work, we focus on separating easily confused classes whose concepts/labels lie in the same leaf group on a hierarchical taxonomy, suggested by expert knowledge (i.e., ChatGPT - Figure 3)
@@ 4 Proposed Method
In the previous section, we noted that increasing overlap in data representations as more tasks arrive is one of the main reasons leading to performance degradation (i.e., forgetting). Therefore, if we can identify easily confused classes/concepts, we will better enhance the distinguishability of models and thus reduce forgetting. Inspired by Cognitive Science studies, regarding the benefits of organizing information in a meaningful way (Section 4.1), we propose using expert knowledge to structure the concepts/labels of CL tasks in the corresponding hierarchical taxonomies. Interestingly, we find that concepts within the same group in each taxonomy tend to be visually and semantically similar, potentially causing more overlap in the latent space and confusion for the CL classifier. Motivated by this observation, we propose group-based contrastive learning in Section 4.2 to maximize the separability of these concepts, thereby mitigating forgetting over time. Additionally, to reinforce the effectiveness of this strategy, Section 4.3 presents an optimal transport-based strategy, which harnesses prior knowledge of pretrained backbones from a new perspective.
@@ 4.1 Motivation
Figure 3: The hierarchical taxonomy obtained when learning Task 3 on Split-CIFAR100.
@@
Figure 4: t-SNE visualization of classes within leaf groups of Four-legged animals ($\bullet$ circular points) and Plants ($\blacktriangle$ triangular points) when learning Task 3, Split-CIFAR100. The order of appearance of the classes: Task 1 - ”mouse”, ”porcupine”, ”oak tree”; Task 2 - ”willow tree”; Task 3 - ”otter”, ”hamster”, ”pine tree” (Figure 2 and 3). We can see that if we train tasks independently with $\mathcal{L}_{CE}$, the classes within each such group, which arrive at different tasks, can be overlapped seriously (Figure 4(a)).
@@
Taking the learning process of Split-CIFAR100 (see Appendix A.1) as an example, when training on task $t=3$, we can construct a tree-like taxonomy of concepts/labels, as shown in Figure 3. We observe that the classes under a same leaf in the structure (e.g., oak tree, willow tree, and pine tree in the leaf group ”plants”) exhibit stronger visual and semantic correlations. Consequently, as shown in Figure 4a, these class features are more overlapping with each other rather than with another class from other leaf groups (e.g., ”four-legged”). In this way, the tree-like taxonomy serves as a tool to help identify easily confused classes, facilitating the subsequent process of making them more separable in the feature space. Leveraging the insight obtained from the taxonomy above, we aim to train a network so that all class representations must be distinct, especially those within each leaf group. Over time, this strategy not only helps the model better mitigate forgetting caused by representation overlap but also reinforces existing knowledge to facilitate future learning.
@@ 4.2 Leveraging Hierarchical Taxonomies for Better Representation Learning
During the training process, whenever a new class appears, its label name is automatically added to the tree-like taxonomy, into a leaf group containing classes with similar characteristics (Figure 3). To develop this hierarchical structure, we can rely on expert knowledge, for example, ChatGPT, which can help us incrementally construct a meaningful and semantically related tree (see Appendix B). Structuring information in this way not only aligns with how the human brain effectively connects and remembers information but also provides useful insights during training, indicating how each knowledge is related to the other and which requires further focus.
For each task $t$, we dedicate a set of prompt $\bm{P}_{t}$. In each task, we aim to minimize overlap between all classes so far, especially focusing on increasing the separability between classes belonging to the same leaf group on the taxonomy (e.g., four-legged mammals, plants, etc.,). That is, taxonomy acts as a reference information channel to support the training process. Let $g$ be a leaf group on the taxonomy $\mathcal{G}$ (i.e., $g\in\mathcal{G}$), ${X}^{g}_{k}$ and ${Y}^{g}_{k}$ denote the corresponding sets of input samples and labels under the group $g$ that belong to the task $k$ ($k\leq t$). Besides Cross Entropy loss $\mathcal{L}_{CE}$, we propose using a regularization loss function for sample ${\bm{x}}$ that arrives in task $t$ and belongs to leaf group $g$ as follows:
@@
$\mathcal{L}_{g}(\psi,\bm{P}_{t},\bm{x})$ $=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
$\displaystyle\mathcal{L}_{all}(\psi,\bm{P}_{t},\bm{x})$ $=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
are the Supervised Contrastive loss that we put on class representation within leaf group $g$, and all classes so far, respectively; $u({\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}})=\exp(\frac{{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}}}{\tau})$, with ${\bm{z}}_{\bm{x}}=f_{\Phi,\bm{P}_{t}}({\bm{x}})$ is the feature vector on the latent space of the prompt-based model and $y_{\bm{x}}$ is the ground truth label of ${\bm{x}}$, $\tau$ is the temperature with $\tau=0.1$ for all experiments, and $\alpha$ is the coefficient that controls how much we want to force classes within leaf group $g$ stay apart further.
In addition, because the data of old tasks can not be accessed when training the new one, we encode the information of each trained class $c$ in the form of a Gaussian Mixture model $\text{GMM}_{c}=\{\mathcal{N}(\mu_{c,i},\Sigma_{c,i})\}_{i=1}^{K}$, at the end of the task, where $K$ is the number of components. Then in task $t$, the representation of each old class $c$ is sampled from $\text{GMM}_{c}$, to compute loss functions in Eq. (1), supporting the training of the current task.
Looking closer at Eq. (1), this equation implies that when learning a new task, the model will be encouraged to identify the decision boundary between old and new classes, especially focusing on those in the same leaf group (see Figure 2). This manipulation matches the way our brain naturally works - without reflection and comparison, we would be confused about old and new concepts, especially those sharing many common characteristics, leading to incorrect judgments and decisions in practice. For A.I models, learning new tasks without thoroughly considering the learned information of old tasks can lead to uncontrolled overlapping of layer representations in the latent space, resulting in forgetting and harming final performance. In this work, thanks to the insight from the hierarchical taxonomy, we can identify which group of classes that easy to get overlapped in the latent space, thereby actively intervening and enhancing the model representation learning.
Figure 5: Wasserterin distance between classes (Split-CIFAR100) in latent space of pretrained backbone (Sup-21K).

### v4

@@ 1 Introduction
Continual Learning (CL) [48, 30] is a research direction that focuses on realizing the human dream of creating truly intelligent systems, where machines can learn on the fly, accumulate knowledge, and operate in constantly changing environments as a human’s companion. Despite the impressive capabilities of A.I systems, continual learning remains a challenging scenario due to the tendency to forget obtained knowledge when facing new ones, known as catastrophic forgetting [15]. In dealing with this challenge, traditional CL methods often rely on storing past data for replaying during new tasks [31, 6], which can raise concerns about memory usage and privacy. Besides, prior work shows that replay methods result in overfitting and poor generalization [31, 45, 9]. To overcome these limitations, recent methods leverage the strong generalization ability of pretrained models [17, 21] to solve sequences of CL tasks. A notable line of work is the prompt-based approach [51, 41, 27], where a small set of learnable prompts is injected to pretrained backbones for adapting emerging tasks over time.
@@
In addition, we discover that human natural learning behavior has many valuable aspects, especially the habit of analyzing data, organizing them in a meaningful way, and finding connections between old and new knowledge [40, 3, 42, 33]. Inspired by these practices, we investigate the characteristics of common benchmark datasets, as well as the behavior of pretrained models. Our findings reveal that the data can always be categorized into consistent groups, regardless of their arrival times. Each group usually includes classes with similar semantic information that models may confuse and, thus, should be paid special attention to during training. Identifying these kinds of groups will benefit in improving model representation learning, as well as aiding in addressing catastrophic forgetting.
Therefore, we propose a novel training strategy that constantly arranges emerging data in groups, following a tree-like taxonomy. In particular, during training a new task, models are trained to distinguish all classes so far, especially focusing on classes within the same leaf group. We observe that images belonging to concepts/labels within each of these leaf groups share strong visual and semantic correlations, leading to overlap in the latent space, which compromises performance. By encouraging models to contrast these classes more distinctly, we can reduce the overlap of easily confused ones. This strategy not only mitigates forgetting when new classes emerge but also consolidates domain-specific knowledge in each leaf group. In addition, based on human learning habits, we find that individuals with stronger foundational knowledge absorb new information faster and more effectively. Thus, we propose an optimal transport-based technique to further utilize priori from pretrained models, where their initial behaviors will provide another perspective on the relationships between classes.
@@
  • Inspired from Cognitive Science, we propose a novel approach to reduce forgetting by examining the relationships between data. By dynamically constructing label-based hierarchical taxonomies and leveraging the prior knowledge of pretrained models through an optimal transport-based approach, we can identify the challenging knowledge areas that require further focus during the sequence of tasks.
  • Experimentally, our method demonstrated significant superiority over state-of-the-art methods on various continual learning benchmarks.
@@ 2 Related Work
Class Incremental Learning (CIL).
This is one of the most challenging and widely studied CL scenarios [44, 47], where task identity is unknown during testing, and data of previous tasks is inaccessible during current training [32, 39, 19, 16]. This work follows the setting of CIL and proposes a novel approach for prompt-based CL models.
@@
This line of work exploits the power of pretrained backbone to quickly adapt to the sequence of downstream tasks by updating just a small number of parameters/ prompts for different tasks. In initial work like [52, 51, 41], the absence of explicit training constraints often leads to feature overlapping between classes from different tasks. Therefore, recent methods employ some types of contrastive loss [47, 26] or utilize Vision Language models [49, 36] to better separate features from tasks. Nevertheless, they treat all classes equally during training, missing the opportunity to efficiently distinguish classes, which have many similarities and are easily confused. In this work, we propose a novel approach to reveal the relationships within data, allowing the model to recognize groups of these classes, and develop a deeper understanding of the respective knowledge areas, thereby effectively reducing forgetting.
Using hierarchical label structures
In Deep Learning, existing work [11, 8, 56, 28] have considered using hierarchical label systems for efficient representation learning. However, these approaches require all labels in advance, which is not suitable for Continual Learning, and the trees are often fixed in both depth and width. Recent researches in CL [25, 7] have considered hierarchical structures to either expand the label set to create new learning scenarios or manage memory buffers for rehearsal settings. However, these structures are still limited in representing information and especially cannot indicate the semantic relationships between labels for effective training. Alternatively, we propose to dynamically build a label-based taxonomy over time for rehearsal-free models, where it is not necessary to provide all labels in advance. Additionally, the tree construction will have the support of experts (i.e., human experts or LLMs), allowing for the acquisition of detailed information and the correlation between class data, thereby enhancing the learning of new knowledge and effectively reducing the forgetting of old information.
@@ 3 Preliminary Analysis
In this part, we first present the problem formulation of prompt-based CL in Section 3.1. Following this, we analyze the sources of forgetting in these methods in Section 3.2.
@@ 3.1 Problem formulation
Dataset Feature shift
Split-Imagenet-R 0
Split-CIFAR-100 0
Split-CUB-200 0
(a) Feature shift corresponding to $\mathcal{D}_{1}$ after learning sequence of tasks.
(b) ”Within task accuracy” on $\mathcal{D}_{1}$.
(c) Model accuracy on $\mathcal{D}_{1}$
Figure 1: Empirical study about forgetting. We setup the experiment to eliminate factors (I) and (II), which result in feature shift after learning the sequence of tasks - Table (a). Therefore, the ”within task accuracy” on $\mathcal{D}_{1}$, using classification head $s_{1}(\bm{x})$ to classify classes within task 1 only, remains over time - Figure (b). However, when using head $h_{\psi}$ to classify all classes observed so far, the model accuracy on $\mathcal{D}_{1}$ decrease significantly - Figure (c), suggesting that besides feature shift, there are other factors that lead to forgetting.
@@ 3.2 Forgetting in Prompt-based Methods
In Continual Learning, forgetting refers to the phenomenon in which performance on previously learned tasks decreases over time. Current prompt-based CL methods, which leverage the power of pretrained models, attribute forgetting either to (I) changes in backbone parameters when using the common prompt pool $\bm{P}$ for all tasks [52, 51] or to (II) the inherent mismatch between the models used at training and testing, as discussed and analyzed in [57, 43]. In this work, we complement these views by providing empirical studies to offer new insights about the reason for forgetting in this type of model, which arises from overlapping between old and new emerging class representations.
@@
To uncover the new factor causing forgetting, we deliberately consider the cases where the behavior of backbone w.r.t each learned task is unchanged over time - meaning (I) and (II) would not happen. Firstly, to eliminate concerns about changing learned parameters (I), we consider methods that propose using a distinct set of prompts $\bm{P}_{t}$ to a specific task $t$. Thus, we conduct experiments on HiDE-Prompt [47], which is the latest SOTA in prompt-based CL. Then, the remaining potential factor is the difference between the prompt chosen at inference time. To this end, the training is carried out as usual; when testing, we intentionally choose the right prompt for each sample to ensure that there is no change in the backbone’s behavior compared to training. And the features after backbone will be classified normally without taskID.
@@
For a detailed examination, we propose an experimental setup to analyze the performance degradation of corresponding models over time. Specifically, we train a classification head $s_{1}(\bm{x})$ on the latent space $f_{\Phi,\bm{P}_{1}}(\bm{x})$ derived from the first task $\mathcal{D}_{1}$. As the model undergoes continual training on a sequence of subsequent tasks, we assess the ability of $s_{1}(\bm{x})$ to classify instances from $\mathcal{D}_{1}$ over time, defining this as within-task accuracy. This is different from model accuracy on $\mathcal{D}_{1}$, which is obtained when we use the usual classification head $h_{\psi}$ as the original design of HiDE-Prompt. This classification head $h_{\psi}$ considers all classes that the model has observed so far, instead of just data of $\mathcal{D}_{1}$ as $s_{1}(\bm{x})$. The results in Figure 1(b), show that within-task accuracy definitely remains. In contrast, Figure 1(c), illustrates a significant decrease in true accuracy, raising the question of whether we have overlooked additional factors contributing to final forgetfulness.
@@
Additionally, considering the inference feature space of $f_{\Phi,\bm{P}}(\bm{x})$, we can see that as more tasks arrive, the number of classes increase, making the space fuller and increasing the possibility of overlap between class distributions. To support this point, we illustrate the representations of some class images in Figure 2, as well as the corresponding t-SNE visualizations in Figure 4. In particular, after Task 1, we have representations of ”oak tree”, ”mouse” and ”porcupine” located in quite separate locations. However, when Task 2 and then Task 3 arrive, the appearance of ”willow tree” and ”pine tree” makes the latent space become fuller, and ”oak tree” no longer maintains the separation from the remaining classes as before, leading to a remarkable drop in performance of previous tasks.
@@
Figure 2: Problem: When new classes arrive, the latent space of a model for all tasks so far becomes fuller, and class representations tend to be overlapped, leading to performance degradation in old tasks. Our solution: We focus on separating easily confused classes whose concepts/labels lie in the same leaf group on a label-based hierarchical taxonomy, suggested by expert knowledge (i.e., ChatGPT - Figure 3). [Best viewed in color mode]
@@ 4 Proposed Method
In the previous section, we noted that increasing overlap in data representations as more tasks arrive is one of the main reasons leading to performance degradation (i.e., forgetting). Therefore, if we can identify easily confused classes/concepts, we will better enhance the distinguishability of models and thus reduce forgetting. Inspired by Cognitive Science studies, regarding the benefits of organizing information in a meaningful way (Section 4.1), we propose using expert knowledge to structure the concepts/labels of CL tasks in the corresponding hierarchical taxonomies. Interestingly, we find that concepts within the same group in each taxonomy tend to be visually and semantically similar, potentially causing more overlap in the visual latent space and causing confusion for the CL classifier. Motivated by this observation, we propose group-based contrastive learning in Section 4.2 to maximize the separability of these concepts, thus mitigating forgetting over time. Furthermore, to reinforce the method’s effectiveness, Section 4.3 presents an optimal transport-based technique, harnessing priori of pretrained backbones from a new perspective.
@@ 4.1 Motivation
Figure 3: The label-based hierarchical taxonomy obtained when learning Task 3 on Split-CIFAR100. The colors (i.e., blue, red, orange) of the label names represent the order in which the corresponding classes appear. Accordingly, the tree-like taxonomy is gradually developed and detailed over time. [Best viewed in color mode]
@@
Figure 4: t-SNE visualizations of classes within leaf groups of Four-legged animals ($\bullet$ circular points) and Plants ($\blacktriangle$ triangular points) when learning Task 3, Split-CIFAR-100. The appearance order of the classes: Task 1 - ”mouse”, ”porcupine”, ”oak tree”; Task 2 - ”willow tree”; Task 3 - ”otter”, ”hamster”, ”pine tree” (also refer to Figure 2 and 3). We can see that if we train tasks independently with $\mathcal{L}_{CE}$, the classes within each leaf group, which arrive at different time, can be overlapped seriously (Figure 4(a)).
@@
Taking the learning process of Split-CIFAR-100 (see Appendix A.1) as an example, when training on task $t=3$, we can construct a tree-like taxonomy of concepts/labels, as shown in Figure 3. We observe that the classes under a same leaf in the structure (e.g., oak tree, willow tree, and pine tree in the leaf group ”plants”) exhibit stronger visual and semantic correlations. Consequently, as shown in Figure 4a, these class features are more overlapping with each other rather than with another class from other leaf groups (e.g., ”four-legged”). In this way, the tree-like taxonomy serves as a tool to help identify easily confused classes, facilitating the subsequent process of making them more separable in the feature space. Leveraging the insight obtained from the taxonomy above, we aim to train a network so that all class representations must be distinct, especially those within each leaf group. Over time, this strategy not only helps the model better mitigate forgetting caused by representation overlap but also reinforces existing knowledge to facilitate future learning.
@@ 4.2 Leveraging Hierarchical Taxonomies for Better Representation Learning
During the training process, whenever a new class appears, its label name is automatically added to the tree-like taxonomy, into a leaf group containing classes with similar characteristics (Figure 3). To develop this hierarchical structure, we can rely on expert knowledge, for example, ChatGPT, which can help incrementally construct a meaningful related tree (see Appendix B). Structuring information in this way not only aligns with how the human brain effectively connects and remembers information but also provides useful insights during training, indicating how each knowledge is related to the other and which requires further focus.
Particularly, for each task $t$, we dedicate a set of prompt $\bm{P}_{t}$. We aim to minimize overlap between all classes so far, especially focusing on increasing the separability between classes belonging to the same leaf group on the taxonomy (e.g., four-legged mammals, plants, etc.,). That is, taxonomy acts as a reference information channel to support the training process. Let $g$ be a leaf group on the taxonomy $\mathcal{G}$ (i.e., $g\in\mathcal{G}$), ${X}^{g}_{k}$ and ${Y}^{g}_{k}$ denote the corresponding sets of input samples and labels under the group $g$ that belong to the task $k$ ($k\leq t$). Besides Cross Entropy loss $\mathcal{L}_{CE}$, we propose using a regularization loss function for sample ${\bm{x}}$ that arrives in task $t$ and belongs to leaf group $g$ as follows:
@@
$\mathcal{L}_{g}(\cdot)$ $=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{1...t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{1...t}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
$\displaystyle\mathcal{L}_{all}(\cdot)$ $=-\log\sum_{{\bm{x}}^{\prime}\in X_{1...t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X_{1...t}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
are the Supervised Contrastive losses that we put on class representation within leaf group $g$, and all classes so far, respectively; and
$u({\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}})=\exp(\frac{{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}}}{\tau})$ (2)
with ${\bm{z}}_{\bm{x}}=f_{\Phi,\bm{P}_{t}}({\bm{x}})$ is the feature vector on the latent space of the prompt-based model and $y_{\bm{x}}$ is the ground truth label of ${\bm{x}}$, $\tau$ is the temperature with $\tau=0.1$ for all experiments, and $\alpha$ is the coefficient that controls how much we want to force classes within leaf group $g$ stay apart further. In addition, because the data of old tasks can not be accessed when training the new one, we follow previous work [47, 27] to encode the information of each trained class $c$ in the form of a Gaussian Mixture model $\text{GMM}_{c}=\{\mathcal{N}(\mu_{c,i},\Sigma_{c,i})\}_{i=1}^{K}$, at the end of the corresponding task, where $K$ is the number of components. Then in task $t$, the representation of each old class $c$ is sampled from $\text{GMM}_{c}$, to compute loss functions in Eq. (1).
Looking closer at Eq. (1), this equation implies that when learning a new task, the model will be encouraged to identify the decision boundary between all old and new classes (i.e., using $\mathcal{L}_{all}$). Besides, it will especially focusing on those in the same leaf group, which often share many common characteristics and can confuse classification models (i.e., using $\mathcal{L}_{g}$), see Figure 2. This manipulation matches the way our brain naturally works - without reflection and comparison, we would be confused about old and new concepts, especially those sharing many common characteristics, leading to incorrect judgments and decisions in practice. For AI models, learning new tasks without thoroughly considering the learned information of old tasks can lead to uncontrolled overlapping in the latent space, resulting in forgetting the learned knowledge and harming final performance. In this work, thanks to the insight from the hierarchical taxonomy, we can identify which group of classes that easy to get overlapped in the latent space, thereby actively intervening and enhancing the model representation learning in the CL environment.